# decoding inner speech with an end-to-end brain-to-text neural interface

### v2

@@ 1 Introduction
Speech neuroprosthetics have experienced a recent revolution, enabling the decoding of attempted and imagined speech and restoring communication for those who have lost the ability to speak (akbari2019towards; makin2020machine; moses2021neuroprosthesis; proix2022imagined; willett2023high; fan2023plug; metzger2023high; card2024accurate; littlejohn2025streaming; kunz2025inner). Despite this success, most existing systems rely on cascaded frameworks that cannot easily be optimized end-to-end. In such models, recurrent neural networks (RNNs) map neural activity to phonemes that are then assembled into sentences with an n-gram LM (willett2023high). While efficient for small datasets, separate optimization of each component can result in a dissociation between the performance of the RNN and the system as a whole. For instance, lower phoneme error rates (PER) from the RNN do not always translate to lower word error rates (WER) when decoding with the n-gram model (willett2024brain). These limitations underscore the need for a fully end-to-end optimized speech decoding framework.
@@
Recently, feng2024towards connect RNNs with LLMs to translate neural activity into sentences in an end-to-end manner. While promising, this approach does not explore modern architectures, such as transformers, which may better capture complex neural representations and enhance decoding performance. In contrast, feghhi2025time incorporate transformers to decode phonemes and apply time masking to mitigate overfitting, achieving higher performance, but their study does not employ end-to-end training with LLMs or leverage large-scale pretraining. Since transformers typically require large datasets to reach their full potential, we hypothesize that combining transformer-based neural encoders with large-scale pretraining could further improve decoding performance and provide stable representations to effectively guide LLMs (liu2023visual).
@@
Here, we present an end-to-end BraIn-to-Text neural interface that we call BIT, which translates neural activity directly into coherent sentences. Our approach introduces several key innovations: a transformer neural encoder pretrained with self-supervised masked modeling on 367 hours of Utah array recordings from humans and monkeys across speech and arm-related motor tasks; and, after fine-tuning, a flexible decoder that supports either a cascaded framework, where decoded phonemes are assembled into sentences with an n-gram LM, or an end-to-end framework, where an audio-LLM decoder directly generates sentences with contrastive alignment between neural and text embeddings. Analogous to LLAVA (liu2023visual), which augments a pretrained LLM with an image encoder as its eyes for visual perception, BIT equips the LLM with a brain to interpret neural activity. Through transformer-based pretraining and end-to-end LLM integration, BIT overcomes the limitations of RNN-only encoders and small-scale datasets, yielding neural representations that generalize across tasks and align speech-related neural activity with the language structure of LLMs.
@@
We evaluate our approach on the Brain-to-Text Benchmark ’24 and ’25 datasets (willett2023high; card2024accurate), which consist of Utah array recordings from two human participants attempting to speak sentences cued on a computer monitor. First, we benchmark the pretrained encoder on attempted speech decoding, demonstrating that it outperforms both RNNs and transformers trained from scratch in the cascaded framework. Next, we evaluate LLM decoders, showing that small-scale audio-LLMs substantially improve end-to-end decoding performance, surpassing the text-based LLM used in prior work (feng2024towards). These gains are even more pronounced when transferring to imagined speech (kunz2025inner), a low-data task with far fewer labeled examples. Interpretability analysis reveals that BIT’s neural embeddings for attempted and imagined speech align through shared semantic structure. These findings establish BIT as a new paradigm for integrating intracortical BCIs with LLMs, enabling neural activity to guide sentence generation without relying on intermediate phoneme or character outputs. Overall, BIT represents a step toward end-to-end speech BCIs that translate neural activity directly into coherent sentences.
@@ 2 Related Work
Early speech BCIs transform neural signals from auditory and motor cortices into intelligible outputs, enabling real-time and high-performance brain-to-text communication (bouchard2013functional; akbari2019towards; makin2020machine; moses2021neuroprosthesis; willett2023high; fan2023plug; card2024accurate). These cascaded systems typically map activity to phonemes or characters before language model composition. Recent advances introduce transformers for long-range dependencies, streaming implementations, and end-to-end decoding with LLMs (feghhi2025time; littlejohn2025streaming; feng2024towards). Finally, inner speech decoding raises the possibility of communication without attempting to move (kunz2025inner). However, most prior work has relied on cascaded systems that cannot be trained end-to-end, while the previous end-to-end decoder, in turn, has relied on RNN encoders without pretraining, not taking advantage of modern model architectures and training approaches. Our work closes this gap with a new end-to-end framework powered by cross-species, cross-task transformer-based pretraining.
@@
Increasing attention has been directed toward large-scale pretraining to learn generalizable neural representations across modalities including fMRI, calcium imaging, spiking activity, EEG, and EMG (scotti2024mindeye2; thomas2022self; lurz2020generalization; azabou2023unified; azabou2024multi; ryoo2025generalizable; ye2023neural; ye2025generalist; zhang2024towards; zhang2025neural; cui2024neuro; kaifosh2025generic). These emerging “foundation models” improve performance while reducing reliance on task-specific data, and have shown success in behavior decoding, motor imagery, visual reconstruction, region and cell-type classification, mental state decoding, and neural activity prediction (thomas2022self; scotti2024mindeye2; cui2024neuro; azabou2024multi; yu2025vivo; wang2025foundation). In spiking data, POSSM (ryoo2025generalizable) has demonstrated cross-species transfer from monkey to human imagined handwriting. Building on this progress, we explore cross-species, cross-task self-supervised pretraining for decoding human speech.
@@
Recent advances in visual language models (VLMs) illustrate how large pretrained models can integrate disparate modalities. Early models such as CoCa (yu2022coca) rely on cross-attention projector for aligning billions of image–text pairs, achieving strong image-captioning performance but at high computational cost. li2023blip2 introduce a more efficient design by freezing pretrained visual and language backbones and training a compact Q-former to align modalities. More recently, instruction-tuned LLMs such as LLaVA (liu2023llava) replace query-based intermediates with direct projection layers mapping visual features into the LLM embedding space. This progression, from cross-attention to Q-formers to simple projections, demonstrates that as LLMs become more capable, modality alignment becomes less compute-intensive. Inspired by these trends, we investigate whether LLMs can similarly generalize to speech-related neural signals.
@@ 3.1 Dataset
The Brain-to-Text Benchmark ’24 and ’25 datasets (willett2023high; card2024accurate) are used for attempted speech decoding. The ’24 dataset includes 12,100 sentences from participant T12 across 25 sessions over four months, with 128-electrode Utah array recordings from ventral motor cortex. The ’25 dataset contains 10,948 sentences from participant T15 across 45 sessions over 20 months, recorded with 256 electrodes in ventral motor cortex. For imagined speech, we use the inner speech dataset (kunz2025inner), comprising 500 and 712 sentences for T12 and T15, recorded with 128 or 256 electrodes while participants internally spoke cued sentences. For all datasets, neural features include thresholded spike counts and spiking-band power (SBP) (nason2020low), binned in 20 ms windows and normalized across days (z-scored) to address Utah array non-stationarity (karpowicz10stabilizing). For pretraining, we use $\sim$ 98 hours of human Utah array recordings (including the above speech BCI datasets) and $\sim$ 269 hours of monkey recordings during motor tasks (willett2021high; willett2023high; card2024accurate; fan2023plug; karpowicz2024few; kunz2025inner; churchland2012neural; chowdhury2020area; even2019structure; ma2023using; perich2018neural). Since SBP is unavailable in some pretraining datasets, we use only thresholded spikes, normalized across days. See Appendix A for details.
@@ 3.2 Architecture
We employ a transformer to learn latent representations from neural activity. The original data has shape $(T,C)$, where $T$ is the number of time bins and $C$ is the number of electrodes. Following feghhi2025time, we group every $T_{\text{patch}}$ time bins into a patch, yielding data of shape $(T/T_{\text{patch}},C\times T_{\text{patch}})$. These patches are passed through a patch embedding module (LayerNorm, Linear, LayerNorm) to produce tokens for the transformer encoder with multi-headed attention. Using time patches rather than individual 20 ms bins aligns the finer temporal resolution of neural recordings with the slower timescale of speech production (30–60 words per minute). In addition, time patch tokens shorten the context length, minimizing information redundancy when feeding neural embeddings to LLMs for short sentence generation.
Each transformer block includes a self-attention layer followed by a feed-forward network (Appendix Q). Relative positional embeddings (RoPE) (su2024roformer) encode temporal information, and a bidirectional attention mask allows each time patch to attend to all others. For self-supervised pretraining, neural activity from each subject is converted into transformer inputs via the patch embedding module, and the transformer outputs are reconstructed to the original neural data through the reversed patch embedding module, which projects latent tokens back to the input dimensionality. For phoneme decoding, transformer outputs are passed through a linear layer to generate logits over phonemes, the blank token, and the silence token.
@@
We apply low-rank adaptation (LoRA) (hu2022lora) to a subset of LLM parameters, including attention projections and feed-forward layers, to enable fine-tuning with few trainable parameters. For audio-based LLMs, LoRA is also applied to the multimodal projector, mapping neural embeddings into the audio embedding space to align with the model’s acoustic representations.
@@ 3.3 Training Objectives
For pretraining, we utilize a temporal masking strategy inspired by masked autoencoders (he2022masked). Neural activity is tokenized into temporal patches, and a subset is randomly replaced with a learnable mask token. Masked patches can form contiguous spans of variable length, up to a predefined maximum timespan, while maintaining a consistent overall masking ratio (Appendix Q). The model is then trained to reconstruct the original neural activity from these partially masked sequences, encouraging the learning of rich contextual representations without external supervision. We use Mean Squared Error (MSE) as the reconstruction loss, since both thresholded spike counts and SBP are normalized. This masking strategy reduces overfitting through data augmentation and mitigates non-stationarity from probe drift (karpowicz10stabilizing; karpowicz2024few). Additionally, pretraining is limited to human and primate Utah array data, allowing the model to learn stable probe representations that are robust to variations in electrode placement, subjects, and behavioral tasks.
@@
After pretraining, the neural encoder is finetuned to predict phoneme sequences from neural activity using a Connectionist Temporal Classification (CTC) loss (graves2006connectionist). The encoder transforms neural activity into latent embeddings, which are projected onto phoneme classes plus a blank token to produce phoneme logits. During fine-tuning, the time masking module is removed, since pretraining with extensive data augmentations has already mitigated overfitting. In the end-to-end model, phoneme logits are not fed to the audio-LLM decoder; rather, they serve as intermediate targets that encode phonetic information into neural representations to guide LLM fine-tuning, even when the goal is sentence prediction (Appendix C).
@@ 4.1 Tasks
We evaluate our approach on two speech decoding tasks. For attempted speech, neural activity is decoded as the participant physically tries to articulate words, although no intelligible audio output is produced due to paralysis (willett2023high; card2024accurate). For imagined speech, neural activity is decoded as the participant silently imagines speaking without attempting to engage the speech muscles (kunz2025inner). Performance is measured by word error rate (WER) averaged across sentences, with lower WER indicating more accurate decoding.
@@ 4.2 Baselines
We first decode phonemes from neural activity using a neural encoder. Phoneme logits are passed into a 5-gram LM to generate sentence hypotheses, re-scored with OPT (zhang2022opt) (Appendix D). Although not fully differentiable, this approach leads to the lowest WER for Utah-array speech BCIs (willett2024brain) and is used to benchmark baseline encoders:
@@
    : Baseline RNN from prior speech BCIs (willett2023high; card2024accurate).
@@
The neural encoder trained for phoneme decoding is paired with a pretrained LLM, fine-tuned to translate neural activity directly into sentences by feeding neural representations, rather than phoneme logits, into the LLM. We benchmark the same neural encoders ( RNN, BIT-TFS, BIT-Human, BIT-All, BIT-Cross-Task-Only ) with various text- and audio-based LLMs (Section 5.2) using nucleus sampling (holtzman2019curious) for sentence generation at inference time (Appendix F). Although end-to-end methods (feng2024towards) have not yet surpassed the cascaded decoder, our approach aims to narrow this gap.
@@ 5.1 Decoding attempted and imagined speech
For attempted speech, we compare BIT-All , BIT-Human , BIT-TFS , and the RNN baseline. For imagined speech, we also include BIT-Cross-Task-Only to evaluate within-subject, cross-task transfer. For both RNN and transformer architectures, we perform hyperparameter tuning by initializing 30 models with optimizer hyperparameters randomly sampled from predefined ranges. We select the configuration with the best validation performance. See Appendix Q-R for details. In our experiments, BIT-TFS outperforms the RNN on attempted speech decoding (Figure 2A); since the RNN does not use time masking, this comparison reflects both architectural and training objective differences. Pretraining at scale ( BIT-Human , BIT-All ) further improves performance over training the encoder from scratch. For imagined speech (50-word vocabulary), BIT-Human and BIT-All outperform BIT-TFS by 39-45% in WER, showing that SSL pretraining is especially beneficial in low-data settings. Moreover, BIT-All exceeds BIT-Cross-Task-Only (Figure 2B–C), indicating that cross-subject, label-free SSL pretraining yields larger transfer gains than same-subject, cross-task supervised pretraining. Although BIT surpasses the prior end-to-end baseline (feng2024towards) combining an RNN encoder with a text-based LLM decoder, a gap in WER remains between cascaded and end-to-end approaches. Our method substantially narrows this gap, bringing end-to-end performance closer to that of the cascaded decoder. Figure 2D shows example sentences decoded by BIT-All in the end-to-end framework.
@@
To benchmark against other methods, we participate in the Brain-to-Text ’24 and ’25 competitions, where WER is measured on a holdout set with ground-truth labels hidden from participants. For fair comparisons, we compare methods in the same decoding framework (cascaded or end-to-end); see Appendix H for details. In Tables 1 and 2, BIT Cascaded (5-gram LM) and BIT End-to-End (Aero1-Audio 1.5B) couple our pretrained encoder ( BIT-All ) with different decoders. BIT Cascaded achieves a SOTA WER of 6.35% for non-ensembled models, surpassing the previous best of 7.98% (feghhi2025time). With a fine-tuned LLM that merges multiple decoded sentences from an ensemble of models trained with different seeds (Appendix G), BIT Cascaded + Ensemble further reduces WER to 5.10%, outperforming the prior 5.68% (feghhi2025time) and ranking first. BIT End-to-End + Ensemble reduces WER from 24.69% to 10.22% compared to the previous end-to-end entry (feng2024towards). On the ’25 leaderboard, BIT Cascaded + Ensemble leads with 2.21% WER, while BIT End-to-End + Ensemble achieves 7.76%.

### v3

# arXiv:2511.21740v3[cs.CL] 28 Feb 2026

@@ 1 Introduction
Speech neuroprosthetics have experienced a recent revolution, enabling the decoding of attempted and imagined speech and restoring communication for those who have lost the ability to speak (Akbari et al., 2019; Makin et al., 2020; Moses et al., 2021; Proix et al., 2022; Willett et al., 2023; Fan et al., 2023; Metzger et al., 2023; Card et al., 2024; Littlejohn et al., 2025; Kunz et al., 2025). Despite this success, most existing systems rely on cascaded frameworks that cannot easily be optimized end-to-end. In such models, recurrent neural networks (RNNs) map neural activity to phonemes that are then assembled into sentences with an n-gram LM (Willett et al., 2023). While efficient for small datasets, separate optimization of each component can result in a dissociation between the performance of the RNN and the system as a whole. For instance, lower phoneme error rates (PER) from the RNN do not always translate to lower word error rates (WER) when decoding with the n-gram model (Willett et al., 2024). These limitations underscore the need for a fully end-to-end optimized speech decoding framework.
@@
Recently, Feng et al. (2024) connect RNNs with LLMs to translate neural activity into sentences in an end-to-end manner. While promising, this approach does not explore modern architectures, such as transformers, which may better capture complex neural representations and enhance decoding performance. In contrast, Feghhi et al. (2025) incorporate transformers to decode phonemes and apply time masking to mitigate overfitting, achieving higher performance, but their study does not employ end-to-end training with LLMs or leverage large-scale pretraining. Since transformers typically require large datasets to reach their full potential, we hypothesize that combining transformer-based neural encoders with large-scale pretraining could further improve decoding performance and provide stable representations to effectively guide LLMs (Liu et al., 2023b).
@@
Here, we present an end-to-end BraIn-to-Text neural interface that we call BIT, which translates neural activity directly into coherent sentences. Our approach introduces several key innovations: a transformer neural encoder pretrained with self-supervised masked modeling on 367 hours of Utah array recordings from humans and monkeys across speech and arm-related motor tasks; and, after fine-tuning, a flexible decoder that supports either a cascaded framework, where decoded phonemes are assembled into sentences with an n-gram LM, or an end-to-end framework, where an audio-LLM decoder directly generates sentences with contrastive alignment between neural and text embeddings. Analogous to LLAVA (Liu et al., 2023b), which augments a pretrained LLM with an image encoder as its eyes for visual perception, BIT equips the LLM with a brain to interpret neural activity. Through transformer-based pretraining and end-to-end LLM integration, BIT overcomes the limitations of RNN-only encoders and small-scale datasets, yielding neural representations that generalize across tasks and align speech-related neural activity with the language structure of LLMs.
@@
We evaluate our approach on the Brain-to-Text Benchmark ’24 and ’25 datasets (Willett et al., 2023; Card et al., 2024), which consist of Utah array recordings from two human participants attempting to speak sentences cued on a computer monitor. First, we benchmark the pretrained encoder on attempted speech decoding, demonstrating that it outperforms both RNNs and transformers trained from scratch in the cascaded framework. Next, we evaluate LLM decoders, showing that small-scale audio-LLMs substantially improve end-to-end decoding performance, surpassing the text-based LLM used in prior work (Feng et al., 2024). These gains are even more pronounced when transferring to imagined speech (Kunz et al., 2025), a low-data task with far fewer labeled examples. Interpretability analysis reveals that BIT’s neural embeddings for attempted and imagined speech align through shared semantic structure. These findings establish BIT as a new paradigm for integrating intracortical BCIs with LLMs, enabling neural activity to guide sentence generation without relying on intermediate phoneme or character outputs. Overall, BIT represents a step toward end-to-end speech BCIs that translate neural activity directly into coherent sentences.
@@ 2 Related Work
Early speech BCIs transform neural signals from auditory and motor cortices into intelligible outputs, enabling real-time and high-performance brain-to-text communication (Bouchard et al., 2013; Akbari et al., 2019; Makin et al., 2020; Moses et al., 2021; Willett et al., 2023; Fan et al., 2023; Card et al., 2024). These cascaded systems typically map activity to phonemes or characters before language model composition. Recent advances introduce transformers for long-range dependencies, streaming implementations, and end-to-end decoding with LLMs (Feghhi et al., 2025; Littlejohn et al., 2025; Feng et al., 2024). Finally, inner speech decoding raises the possibility of communication without attempting to move (Kunz et al., 2025). However, most prior work has relied on cascaded systems that cannot be trained end-to-end, while the previous end-to-end decoder, in turn, has relied on RNN encoders without pretraining, not taking advantage of modern model architectures and training approaches. Our work closes this gap with a new end-to-end framework powered by cross-species, cross-task transformer-based pretraining.
@@
Increasing attention has been directed toward large-scale pretraining to learn generalizable neural representations across modalities including fMRI, calcium imaging, spiking activity, EEG, and EMG (Scotti et al., 2024; Thomas et al., 2022; Lurz et al., 2020; Azabou et al., 2023; 2024; Ryoo et al., 2025; Ye et al., 2023; 2025; Zhang et al., 2024; 2025; Cui et al., 2024; Kaifosh and Reardon, 2025). These emerging “foundation models” improve performance while reducing reliance on task-specific data, and have shown success in behavior decoding, motor imagery, visual reconstruction, region and cell-type classification, mental state decoding, and neural activity prediction (Thomas et al., 2022; Scotti et al., 2024; Cui et al., 2024; Azabou et al., 2024; Yu et al., 2025; Wang et al., 2025). In spiking data, POSSM (Ryoo et al., 2025) has demonstrated cross-species transfer from monkey to human imagined handwriting. Building on this progress, we explore cross-species, cross-task self-supervised pretraining for decoding human speech.
@@
Recent advances in visual language models (VLMs) illustrate how large pretrained models can integrate disparate modalities. Early models such as CoCa (Yu et al., 2022) rely on cross-attention projector for aligning billions of image–text pairs, achieving strong image-captioning performance but at high computational cost. Li et al. (2023) introduce a more efficient design by freezing pretrained visual and language backbones and training a compact Q-former to align modalities. More recently, instruction-tuned LLMs such as LLaVA (Liu et al., 2023a) replace query-based intermediates with direct projection layers mapping visual features into the LLM embedding space. This progression, from cross-attention to Q-formers to simple projections, demonstrates that as LLMs become more capable, modality alignment becomes less compute-intensive. Inspired by these trends, we investigate whether LLMs can similarly generalize to speech-related neural signals.
@@ 3.1 Dataset
The Brain-to-Text Benchmark ’24 and ’25 datasets (Willett et al., 2023; Card et al., 2024) are used for attempted speech decoding. The ’24 dataset includes 12,100 sentences from participant T12 across 25 sessions over four months, with 128-electrode Utah array recordings from ventral motor cortex. The ’25 dataset contains 10,948 sentences from participant T15 across 45 sessions over 20 months, recorded with 256 electrodes in ventral motor cortex. For imagined speech, we use the inner speech dataset (Kunz et al., 2025), comprising 500 and 712 sentences for T12 and T15, recorded with 128 or 256 electrodes while participants internally spoke cued sentences. For all datasets, neural features include thresholded spike counts and spiking-band power (SBP) (Nason et al., 2020), binned in 20 ms windows and normalized across days (z-scored) to address Utah array non-stationarity (Karpowicz et al., ). For pretraining, we use $\sim$ 98 hours of human Utah array recordings (including the above speech BCI datasets) and $\sim$ 269 hours of monkey recordings during motor tasks (Willett et al., 2021; 2023; Card et al., 2024; Fan et al., 2023; Karpowicz et al., 2024; Kunz et al., 2025; Churchland et al., 2012; Chowdhury et al., 2020; Even-Chen et al., 2019; Ma et al., 2023; Perich et al., 2018). Since SBP is unavailable in some pretraining datasets, we use only thresholded spikes, normalized across days. See Appendix A for details.
@@ 3.2 Architecture
We employ a transformer to learn latent representations from neural activity. The original data has shape $(T,C)$, where $T$ is the number of time bins and $C$ is the number of electrodes. Following Feghhi et al. (2025), we group every $T_{\text{patch}}$ time bins into a patch, yielding data of shape $(T/T_{\text{patch}},C\times T_{\text{patch}})$. These patches are passed through a patch embedding module (LayerNorm, Linear, LayerNorm) to produce tokens for the transformer encoder with multi-headed attention. Using time patches rather than individual 20 ms bins aligns the finer temporal resolution of neural recordings with the slower timescale of speech production (30–60 words per minute). In addition, time patch tokens shorten the context length, minimizing information redundancy when feeding neural embeddings to LLMs for short sentence generation.
Each transformer block includes a self-attention layer followed by a feed-forward network (Appendix Q). Relative positional embeddings (RoPE) (Su et al., 2024) encode temporal information, and a bidirectional attention mask allows each time patch to attend to all others. For self-supervised pretraining, neural activity from each subject is converted into transformer inputs via the patch embedding module, and the transformer outputs are reconstructed to the original neural data through the reversed patch embedding module, which projects latent tokens back to the input dimensionality. For phoneme decoding, transformer outputs are passed through a linear layer to generate logits over phonemes, the blank token, and the silence token.
@@
We apply low-rank adaptation (LoRA) (Hu et al., 2022) to a subset of LLM parameters, including attention projections and feed-forward layers, to enable fine-tuning with few trainable parameters. For audio-based LLMs, LoRA is also applied to the multimodal projector, mapping neural embeddings into the audio embedding space to align with the model’s acoustic representations.
@@ 3.3 Training Objectives
For pretraining, we utilize a temporal masking strategy inspired by masked autoencoders (He et al., 2022). Neural activity is tokenized into temporal patches, and a subset is randomly replaced with a learnable mask token. Masked patches can form contiguous spans of variable length, up to a predefined maximum timespan, while maintaining a consistent overall masking ratio (Appendix Q). The model is then trained to reconstruct the original neural activity from these partially masked sequences, encouraging the learning of rich contextual representations without external supervision. We use Mean Squared Error (MSE) as the reconstruction loss, since both thresholded spike counts and SBP are normalized. This masking strategy reduces overfitting through data augmentation and mitigates non-stationarity from probe drift (Karpowicz et al., ; Karpowicz et al., 2024). Additionally, pretraining is limited to human and primate Utah array data, allowing the model to learn stable probe representations that are robust to variations in electrode placement, subjects, and behavioral tasks.
@@
After pretraining, the neural encoder is finetuned to predict phoneme sequences from neural activity using a Connectionist Temporal Classification (CTC) loss (Graves et al., 2006). The encoder transforms neural activity into latent embeddings, which are projected onto phoneme classes plus a blank token to produce phoneme logits. During fine-tuning, the time masking module is removed, since pretraining with extensive data augmentations has already mitigated overfitting. In the end-to-end model, phoneme logits are not fed to the audio-LLM decoder; rather, they serve as intermediate targets that encode phonetic information into neural representations to guide LLM fine-tuning, even when the goal is sentence prediction (Appendix C).
@@ 4.1 Tasks
We evaluate our approach on two speech decoding tasks. For attempted speech, neural activity is decoded as the participant physically tries to articulate words, although no intelligible audio output is produced due to paralysis (Willett et al., 2023; Card et al., 2024). For imagined speech, neural activity is decoded as the participant silently imagines speaking without attempting to engage the speech muscles (Kunz et al., 2025). Performance is measured by word error rate (WER) averaged across sentences, with lower WER indicating more accurate decoding.
@@ 4.2 Baselines
We first decode phonemes from neural activity using a neural encoder. Phoneme logits are passed into a 5-gram LM to generate sentence hypotheses, re-scored with OPT (Zhang et al., 2022) (Appendix D). Although not fully differentiable, this approach leads to the lowest WER for Utah-array speech BCIs (Willett et al., 2024) and is used to benchmark baseline encoders:
@@
    : Baseline RNN from prior speech BCIs (Willett et al., 2023; Card et al., 2024).
@@
The neural encoder trained for phoneme decoding is paired with a pretrained LLM, fine-tuned to translate neural activity directly into sentences by feeding neural representations, rather than phoneme logits, into the LLM. We benchmark the same neural encoders ( RNN, BIT-TFS, BIT-Human, BIT-All, BIT-Cross-Task-Only ) with various text- and audio-based LLMs (Section 5.2) using nucleus sampling (Holtzman et al., 2020) for sentence generation at inference time (Appendix F). Although end-to-end methods (Feng et al., 2024) have not yet surpassed the cascaded decoder, our approach aims to narrow this gap.
@@ 5.1 Decoding attempted and imagined speech
For attempted speech, we compare BIT-All , BIT-Human , BIT-TFS , and the RNN baseline. For imagined speech, we also include BIT-Cross-Task-Only to evaluate within-subject, cross-task transfer. For both RNN and transformer architectures, we perform hyperparameter tuning by initializing 30 models with optimizer hyperparameters randomly sampled from predefined ranges. We select the configuration with the best validation performance. See Appendix Q-R for details. In our experiments, BIT-TFS outperforms the RNN on attempted speech decoding (Figure 2A); since the RNN does not use time masking, this comparison reflects both architectural and training objective differences. Pretraining at scale ( BIT-Human , BIT-All ) further improves performance over training the encoder from scratch. For imagined speech (50-word vocabulary), BIT-Human and BIT-All outperform BIT-TFS by 39-45% in WER, showing that SSL pretraining is especially beneficial in low-data settings. Moreover, BIT-All exceeds BIT-Cross-Task-Only (Figure 2B–C), indicating that cross-subject, label-free SSL pretraining yields larger transfer gains than same-subject, cross-task supervised pretraining. Although BIT surpasses the prior end-to-end baseline (Feng et al., 2024) combining an RNN encoder with a text-based LLM decoder, a gap in WER remains between cascaded and end-to-end approaches. Our method substantially narrows this gap, bringing end-to-end performance closer to that of the cascaded decoder. Figure 2D shows example sentences decoded by BIT-All in the end-to-end framework.
@@
To benchmark against other methods, we participate in the Brain-to-Text ’24 and ’25 competitions, where WER is measured on a holdout set with ground-truth labels hidden from participants. For fair comparisons, we compare methods in the same decoding framework (cascaded or end-to-end); see Appendix H for details. In Tables 1 and 2, BIT Cascaded (5-gram LM) and BIT End-to-End (Aero1-Audio 1.5B) couple our pretrained encoder ( BIT-All ) with different decoders. BIT Cascaded achieves a SOTA WER of 6.35% for non-ensembled models, surpassing the previous best of 7.98% (Feghhi et al., 2025). With a fine-tuned LLM that merges multiple decoded sentences from an ensemble of models trained with different seeds (Appendix G), BIT Cascaded + Ensemble further reduces WER to 5.10%, outperforming the prior 5.68% (Feghhi et al., 2025) and ranking first. BIT End-to-End + Ensemble reduces WER from 24.69% to 10.22% compared to the previous end-to-end entry (Feng et al., 2024). On the ’25 public leaderboard, BIT Cascaded + Ensemble leads with 1.76% WER, while BIT End-to-End + Ensemble achieves 7.76%.